# TwT: Thinking without Tokens by Habitual Reasoning Distillation with Multi-Teachers’ Guidance

### v1

@@ Abstract
Jingxian Xu1thanks:   Work during internship at Microsoft., Mengyu Zhou4thanks:   Corresponding author (mezho@microsoft.com)., Weichang Liu211footnotemark: 1, Hanbing Liu311footnotemark: 1, Shi Han4, Dongmei Zhang4 1Nankai University 2BeijingJiaoTong University 3Tsinghua University 4Microsoft Research
@@ 3.1.1 Quality Selection
where $w_{j}$ denotes the weight assigned to the $j$-th factor and $m_{j}$ represents the corresponding factor value. For example, in a code generation task, we let the large model assign a weighted score based on factors such as the reasoning process, code readability, and robustness, yielding the final confidence score. For $k$ teacher models, we can obtain a set of outputs $O_{i}=\{O_{i1},O_{i2},\dots,O_{ik}\}$ for each query $x_{i}$. Then, we set a confidence score threshold $s$. For each output $O_{i}$, if its confidence score $c_{i}\geq s$, the output sample is considered high-quality and retained; otherwise, it is discarded. Subsequently, diversity selection is performed on the retained sample set =$\mathcal{H}=\{(x_{i},r_{i},y_{i},c_{i})\}_{i=1}^{P}$, where $P$ is the total number of retained samples.
@@ 3.1.2 Diversity Selection
$(p^{*},q^{*})=\arg\min_{1\leq p<q\leq k}\text{sim}\bigl{(}r_{ip},r_{iq}\bigr{)})$ (3)
@@ 3.2 Habitual Reasoning Distillation
In this section, we propose a novel distillation strategy to address the trade-off between inference-time computational cost and model performance. Our approach decomposes the distillation process into three sequential stages. In Stage-1, the student model is distilled using data with full reasoning, allowing it to fully absorb the teacher’s comprehensive thinking process. During the Stage-2, the model learns from data with compressed reasoning, encouraging it to derive solutions using more concise and efficient explanations. Finally, in Stage-3, distilling is conducted on data without any explicit reasoning, enabling the student to learn an end-to-end mapping from query to answer. Notably, in the second stage, we propose a Teacher-Guided Compression method that tailors the complexity of the reasoning information to the student model’s capacity.
@@
where $f$ indicates the student model, $\mathbb{E}_{\mathcal{G}}$ is the expectation over the distillation dataset $\mathcal{G}$, and $\mathrm{P}_{f}([x;r;\hat{y}])$ is the probability assigned by the student model f to the joint input $[x;r;\hat{y}]$.
@@
Stage-3: Reasoning-Free Distillation.
In summary, Stage-1 and Stage-2 use full resoning distillation and reasoning-compressed distillation, respectively, to gradually familiarize the student model with the complete reasoning chain and establish a systematic reasoning pattern. This enables the model to master the key problem-solving steps at both conceptual and operational levels, ultimately internalizing complex reasoning abilities and forming a stable reasoning mechanism. In Stage-3, we perform end-to-end training to further deepen the model’s understanding of the relationship between problems and their answers. Since the explicit reasoning chain has been embedded as an inherent capability in the previous stages, Stage-3 no longer relies on explicit reasoning paths but instead directly reinforces answer generation. This progressive strategy allows the student model to output correct answers without explicit reasoning, thereby reducing inference-time computational cost while enhancing overall performance.
4 Experiment
@@ 4.1 Experiment Setup
Models. We utilize GPT-4, GPT-4omini and Mistral-Large as our teacher models, which are accessed through OpenAI’s API and MistralAI’s API. For the student models, we use Mistral-7B-v0.3 and Phi-3.5mini. For the pre-trained sentence embedding model, we leverage all-mpnet-base-v2.
@@ 4.2 Baseline Comparison
Log Probability-Based 74.83% 75.29% 72.19%
Hard Rejection Sampling-Based 72.29% 73.11% 70.26%
Confidence Score-Based 83.49% 84.90% 81.14%
@@
4.3 Distillation Stage Analysis
4.4 Sampling and Compression Analysis
We evaluated the effectiveness of our DCRS method’s confidence score selection for distillation data and our Teacher-Guided Compression approach for enhancing student model’s performance. For sampling, we compared Confidence Score-Based sampling with two alternatives—Log Probability-Based selection and Hard Rejection Sampling, which categorizes LLM outputs into Good, Medium, and Poor—across three datasets. As shown in Table 4, our method consistently outperforms the alternatives with nearly a 10% gain in accuracy improvement. This highlights the effectiveness of our strategy in identifying high-quality examples for distillation. For compression, we compared our Teacher-Guided Compression with Fixed-Length Compression and a Compressor-based approach. As detailed in Table 5, our method better aligns the teacher’s outputs with the student model’s limited capacity, enhancing both training efficiency and overall performance.
@@ 4.5 Ablation Study
When comparing TwT with the multi-teacher strategy against the single-teacher approach, we observe an average improvement of 1.4% in the distillation performance. This result highlights the effectiveness of the multi-teacher strategy, as it provides a more diverse range of reasoning paths for the student model, thereby facilitating its learning process. Comparing TwT’s DCRS strategy to directly organizing the raw pseudo-labeled data genereated by teachers, TwT achieves an average improvement of 1.5%, underscoring the importance of performing high-quality and diverse data sampling before distillation. This step mitigates the influence of low-quality data and boosts overall distillation performance. Finally, when comparing TwT’s multi-stage distillation approach with the strategy of simply removing the reasoning process (i.e., without any intermediate compression), TwT achieves an approximate 3.5% accuracy gain, while the uncompressed approach not only underperforms in accuracy but also increases the number of output tokens. This finding confirms the necessity of gradually internalizing the reasoning capability for optimal performance.
@@ Appendix A Prompt
A.1 Prompt for Teachers to Generate Pseudo-Labels on the MBPP, CQA and MetaMath Dataset
@@ B.1 Query
A specific query example of the MBPP dataset is shown below:
@@
B.2 Teacher’s Original Response (HaRD Stage-1’s dataset)
After Dual-Criteria rejection sampling, one of the selected teacher responses (with full reasoning process) is:
@@ B.2 Teacher’s Original Response (HaRD Stage-1’s dataset)
B.3 Student’s Inference for The Same Query
After performing SFT on the student model using the full reasoning dataset from the first stage, an initial student model with reasoning ability is obtained. Let this stage’s student model perform inference on the distillation dataset, resulting in the student’s response, as shown in Fig.7
@@ B.4 Teacher’s Refinement Prompt
After the student model performs inference, the reasoning processes of both the student and teacher models are extracted for the same query. The teacher model modifies its original reasoning process based on the student model’s response preferences. The prompt is as follows:
@@
B.5 Compressed Reasoning Data After Refinement (HaRD Stage-2’s dataset)
By having the teacher model refine its original response based on the student model’s answer to the same question, the teacher’s response can better align with the student’s learning characteristics. This can be compared to the human learning process: the teacher delivers the initial lesson, then, after receiving feedback from the student, modifies the lesson plan based on the feedback to improve subsequent teaching. The teacher’s revised response is shown in the Fig.9. Since the refined reasoning is intuitively much shorter than the original reasoning, we refer to it as the “reasoning-compressed dataset”, which is used for the second stage of distillation in HaRD.
Figure 9: Reasoning-compressed distillation stage data.
@@ B.6 Reasoning-free Distillation Data
Appendix C Example Appendix
This is an appendix.

### v2

# Shi Han4, Dongmei Zhang4

@@ Abstract
Jingxian Xu1thanks: Work during internship at Microsoft., Mengyu Zhou4thanks: Corresponding author (mengyu.chou@gmail.com)., Weichang Liu211footnotemark: 1, Hanbing Liu311footnotemark: 1, Shi Han4, Dongmei Zhang4 1Nankai University 2BeijingJiaoTong University 3Tsinghua University 4Microsoft Research
@@ 3.1.1 Quality Selection
where $w_{j}$ denotes the weight assigned to the $j$-th factor and $m_{j}$ represents the corresponding factor value. For example, in a code generation task, we let the large model assign a weighted score based on factors such as the reasoning process, code readability, and robustness, yielding the final confidence score. For $k$ teacher models, we can obtain a set of outputs $O_{i}=\{O_{i1},O_{i2},\dots,O_{ik}\}$ for each query $x_{i}$. Then, we set a confidence score threshold $s$. For each output $O_{i}$, if its confidence score $c_{i}\geq s$, the output sample is considered high-quality and retained; otherwise, it is discarded. Subsequently, diversity selection is performed on the retained sample set $\mathcal{H}=\{(x_{i},r_{i},y_{i},c_{i})\}_{i=1}^{P}$, where $P$ is the total number of retained samples.
@@ 3.1.2 Diversity Selection
$(p^{*},q^{*})=\arg\min_{1\leq p<q\leq k}\text{sim}\bigl(r_{ip},r_{iq}\bigr)$ (3)
@@ 3.2 Habitual Reasoning Distillation
In this section, we introduces a novel multi-stage distillation strategy designed to balance inference-time efficiency and model performance. The process is structured into three sequential stages. In Stage-1, the student model is trained on data with full reasoning, enabling it to internalize the teacher’s comprehensive problem-solving process. In Stage-2, the student learns from compressed reasoning, fostering more concise and efficient inference. In Stage-3, distillation is performed using answer-only data, encouraging the student to establish a direct query-to-answer mapping. Crucially, Stage-2 incorporates a Teacher-Guided Compression mechanism that adaptively aligns the reasoning complexity with the student’s capacity.
@@
[…]
@@
Stage-3: Reasoning-Free Distillation.
In summary, Stage-1 uses full reasoning distillation, while Stage-2 employs compressed reasoning distillation, enabling the student model to learn complete reasoning chains and establish systematic reasoning patterns. Stage-3 conducts end-to-end training to strengthen the model’s understanding of problem-answer relationships, with explicit reasoning now internalized. This progressive approach enhances performance while reducing inference-time computation. The detailed prompt and case study are provided in Appendix A and B.
4 Experiment
@@
Method MBPP CQA MetaMath
TwT-stage1 54.56% 306 64.10% 297 72.91% 155 72.49% 141 45.09% 299 79.17% 320
TwT-stage2 55.31% 162 65.93% 188 75.22% 72 75.62% 74 46.29% 176 80.62% 187
TwT-stage3 56.64% 57 66.16% 113 75.98% 12 76.15% 15 46.93% 16 81.37% 18
Table 4: Quantitative results for three distillation stages via GPT-4, GPT-4o-mini and GPT-3.5-turbo. Accuracy and the number of output tokens were used to evaluate the model performance. The best results were highlighted in bold.
@@ 4.1 Experiment Setup
Models. We utilize GPT-4, GPT-4o-mini and Mistral-Large as our teacher models, which are accessed through OpenAI’s API and MistralAI’s API. For the student models, we use Mistral-7B-v0.3 and Phi-3.5mini. For the pre-trained sentence embedding model, we leverage all-mpnet-base-v2.
@@ 4.3 Analysis on Distillation Stage
Log Probability 74.83% 75.29% 72.19%
Hard Rejection Sampling 72.29% 73.11% 70.26%
Confidence Score 83.49% 84.90% 81.14%
@@
4.4 Analysis on the Robustness to Teacher Model Quality
To evaluate the robustness of our method to the quality of teacher models, we conduct the experiment using teachers with varying capabilities. Specifically, we compare TwT performance when distilled from GPT-4, GPT-4o-mini, and GPT-3.5-turbo. Notably, GPT-3.5-turbo underperforms compared to Mistral-large on our benchmark datasets, particularly on the MetaMath task. Compared Table 2 and Table 4, it demonstrates that TwT remains effective even when guided by a weaker teacher. Despite the reduced capability of GPT-3.5-turbo, the resulting student models still follow our core hypothesis and achieve competitive performance. These findings suggest that TwT is robust to teacher quality and does not rely on the strongest teacher models to be effective.
4.5 Analysis on Sampling and Compression Strategy
We evaluated the effectiveness of our DCRS method’s confidence score selection for distillation data and our Teacher-Guided Compression approach for enhancing student model’s performance. For sampling, Table 5 shows that our Confidence Score-Based method outperformed Log Probability-Based and Hard Rejection Sampling approaches by nearly 10% in accuracy, demonstrating superior quality selection for distillation. For compression, Table 6 shows that Teacher-Guided Compression better matched teacher outputs to student capacity than Fixed-Length or Compressor-based methods, improving both efficiency and performance. Confidence score details are in Appendix C.
@@ 4.6 Ablation Study
TwT’s multi-teacher strategy outperforms single-teacher distillation by 1.4%, demonstrating the benefit of diverse reasoning paths. Compared to using raw pseudo-labeled data directly, TwT’s DCRS strategy shows a 1.5% improvement, highlighting the value of quality data sampling. Additionally, TwT’s multi-stage distillation achieves 3.5% higher accuracy than uncompressed reasoning while reducing output tokens, proving the effectiveness of gradual reasoning internalization.
@@ Appendix A Prompt
The specific prompt for teachers to generate pseudo-labels on the MBPP, CQA and MetaMath dataset is shown in Figure 4.
@@ B.1 Query
A specific query example of the MBPP dataset is shown in Figure 5.
@@
B.2 Teacher’s Original Response (Dataset of HaRD Stage-1)
After Dual-Criteria rejection sampling, one of the selected teacher responses (with full reasoning process) is shown in Figure 6.
@@ B.2 Teacher’s Original Response (Dataset of HaRD Stage-1)
B.3 Student’s Inference for Same Query
After performing SFT on the student model using the full reasoning dataset from the first stage, an initial student model with reasoning ability is obtained. Let this stage’s student model perform inference on the distillation dataset, resulting in the student’s response, as shown in Figure 7.
@@ B.4 Teacher’s Refinement Prompt
After the student model performs inference, the reasoning processes of both the student and teacher models are extracted for the same query. The teacher model modifies its original reasoning process based on the student model’s response preferences. The prompt is shown in Figure 8.
@@
B.5 Compressed Reasoning Data after Refinement (Dataset of HaRD Stage-2)
By having the teacher model refine its original response based on the student model’s answer to the same question, the teacher’s response can better align with the student’s learning characteristics. This can be compared to the human learning process: the teacher delivers the initial lesson, then, after receiving feedback from the student, modifies the lesson plan based on the feedback to improve subsequent teaching. The teacher’s revised response is shown in Figure 9. Since the refined reasoning is intuitively much shorter than the original reasoning, we refer to it as the “reasoning-compressed dataset”, which is used for the second stage of distillation in HaRD.
Figure 9: Dataset of reasoning-compressed distillation stage.
@@ B.6 Reasoning-free Distillation Data
Appendix C Details of Confidence Score Selection
Taking the MBPP dataset as an example, we prompt the LLM to assign a confidence score to its own output based on four weighted criteria: correctness (50%), readability (20%), execution efficiency (20%), and test coverage (10%). The final score is returned in the format score: 0.85 as a clear numeric value.
Inspired by recent work on LLM self-consistency, which shows that LLMs can assess the quality of their own outputs by assigning confidence scores, we adopt a similar strategy to evaluate the trustworthiness of generated samples. Specifically, we introduce a confidence threshold: samples with scores above this threshold are retained as reliable, while those below are discarded.
To determine this threshold, we analyze the relationship between confidence scores and output quality. On datasets such as MBPP, CQA, and MetaMath, we observe that when the confidence score is around 0.95, the alignment between model predictions and ground truth is the highest (e.g., the largest number of passed test cases on MBPP). Based on this observation, we set the confidence threshold to 0.95 in our experiments. The confidence score distributions on different datasets are shown in Figures 1012.
Figure 10: Confidence Score Distribution on MBPP dataset via GPT-4, GPT-4o-mini and Mistal-Large.
Figure 11: Confidence Score Distribution on CQA dataset via GPT-4, GPT-4o-mini and Mistal-Large.
Figure 12: Confidence Score Distribution on MetaMath dataset via GPT-4, GPT-4o-mini and Mistal-Large.